OpenStudy (anonymous):

precal :-/ An airplane is sighted at the same time by two ground observers who are 3 miles apart and both directly west of the airplane. They report the angles of elevation as 13 degrees and 20 degrees. How high is the airplane?

OpenStudy (anonymous):

Make two right triangles, one with base =x and the other with base= 3+x. The nearer one has the 20o angle to the plane. The farther one has the 13o angle height = x tan (20o)=.364 x height = (x + 3) tan (13o) = .231 (x+3) set height = height to get x substitute x into one of the equations to get height.